OpenStudy (anonymous):

If f(x) = -x^2-2x-1, find f(-3)

OpenStudy (anonymous):

why would it be a negative -9? isn't -x going to make it -(-3)^2-2(-3)-1

OpenStudy (lalaly):

f(-3)=-(-3)^2-2(-3)-1= -9+6-1= -3-1=-4 :D

OpenStudy (anonymous):

Hmm.. still a tad confused isn't a negative x negative = positive? So wouldn't -1(-3)^2 = 9

jimthompson5910 (jim_thompson5910):

no, (-3)^2 = 9 but -(-3)^2 = -(9) = -9

OpenStudy (anonymous):

ahh ok I get it now, cause of the square it's three -'s

OpenStudy (anonymous):

Wow little stuff destroys, lol.